OPTICAL ABSORPTION SENSITIVITY BETTER THAN 1 × 10^{−12}

JUN YE, Quantum Optics Group, CalTech, Pasadena CA 91125; MA LONG-SHENG, Quantum Physics Lab, East China Normal University, Shanghai, PR China; \underline{JOHN L. HALL}, and MATTHEW S. TAUBMAN, JILA, University of Colorado and NIST, Boulder CO 80309-0440.

Improved physical understanding of the REAL sensitivity-limiting processes, along with better technical solutions for cavity-enhanced optical heterodyne spectroscopy recently have allowed remarkable sensitivity improvements, presently at the absorption level of 5.2 x10^{−13} for a 1 s integration ^{a}. Here we review the several problems which led to this NICE-OHMS solution, and report recent progress with active control of the Residual Amplitude Modulation produced by the Electro-Optic Modulator. Also, saturated absorption spectra near 1064 nm for HCCD, HCCH, and CO_{2} are presented. The two additional lines are ^{12}C_{2}H_{2} (2ν_{1} + ν_{2} + ν_{5}) R(12) ^{b} and ^{12}C^{16}O_{2} (2ν_{1} + 3ν_{3}) R(6) ^{c}, with their respective transition dipole moments of 50 μDebye and 6 μDebye. They are both weaker than our usual C_{2}HD(ν_{2} + 3 ν_{3}) P(5) transition, which has a transition dipole moment of 70 μDebye, but all are recovered with excellent signal-to-noise ratios. The absolute resonance center frequencies of all three transitions have been measured (+/- 25 kHz) using as reference a Nd:YAG laser locked via frequency doubling on the a_{10} hyperfine-structure component of the R(56) 32-0 I_{2} transition. The C_{2}H_{2} resonance is about 4-fold weaker than that of C_{2}HD, while the pressure broadening rate of 34(1) MHz/Torr (FWHM) is similar. For the CO_{2} transition, however, the saturated absorption signal is much weaker, by more than a factor of 350, and shows an elegant and unexpected lineshape which is believed to result from nearly overlapping one- and two-photon transitions.

aJun Ye, Long-Sheng Ma, and John L. Hall, J. Opt. Soc. Am. B15, 6 (Jan 1998).

bK. Nakagawa, T. Katsuda, M. de Labachelerie, and M. Ohtsu, Opt. Comm. 107, 369 (1994).

cP. Fritschel and R. Weiss, Appl. Opt. 31,1910 (1994).